Accounts Receivable Turnover and Number of Days' Sales in Receivables

Financial statement data for years ending December 31 for Blum Company are shown below.

2016 2015
Sales $5,794,740 $5,432,660
Accounts receivable:
  Beginning of year 423,110 474,240
  End of year 496,690 416,360

a. Determine the accounts receivable turnover for 2016 and 2015. Round your answers to one decimal place.

Accounts Receivable Turnover
2016
2015

b. Determine the number of days' sales in receivables for 2016 and 2015. Use 365 days in a year in your computations. Round interim calculations and your final answers to one decimal place.

Number of Days' Sales in Receivables
2016 days
2015 days

c. Does the change in accounts receivable turnover and the number of days' sales in receivables from 2015 to 2016 indicate a favorable or an unfavorable trend?
Answer #1

a) Accounts Receivable Turnover

Accounts Receivable Turnover measures the company's ability to convert its debtors into cash.

Accounts Receivable Turnover (ART) = Net Credit Sales / Average accounts receivables

ART (2015) = $54,32,660 / ($474,240 + $416,360)/2

= 12.2

ART (2016) = $ 57,94,740 / ($423,110 + $ 496,690)/2

= 12.6

b) No.of days sales in receivables

Days sales outstanding is the no.of days the company has taken to convert its accounts receivable.

DSO = 365 days / Accounts Receivable Turnover

DSO (2015) = 365 / 12.2 = 29.92 days

DSO (2016) = 365 / 12.6 = 28.97 days

c) The accounts receivable turnover has improved in 2016 compared to 2015. The days sales in receivables has been reduced. This indicates that the receivables are collected faster than the previous year. Hence, there is an improvement and it indicates a favourable trend.